8. Purchasing-power parity

Using data from The Economist's Big Mac Index for 2019, the following table shows the local currency price of a Big Mac in several countries as well as the actual exchange rate between each country and the United States. At the time of the data collection, a Big Mac would have cost you $5.74 in the United States and GBP 3.29 in the United Kingdom. The actual exchange rate between the British pound and the U.S. dollar was $1.25 per pound. The dollar price of a Big Mac purchased in the United Kingdom was, therefore, computed as follows: 

$$ \begin{aligned} &\text { Dollar price of a Big Mac in the United Kingdom }=\text { GBP } 3.29 \times \frac{\$ 1.25}{\text { GBP } 1.00}\\ &=\$ 4.11 \end{aligned} $$


For the price you paid for a Big Mac in the United States, you could have purchased a Big Mac in the United Kingdom and had some change left over for fries! 

Complete the final column of the table by computing the dollar price of a Big Mac for the countries where this amount is not given. Note: Round your answers to the nearest cent.
Answer #1

$ price in Euro area = 4.08*0.30 = 4.5696

$ price in Norway = 42*0.12 = 5.04

the exchange rate that would have equalized the $ price of Big Mac in the U.S and Euro area(i.e the PPP Exchange for Big Macs) is 5.74/4.08 = 1.40 per euro. this change would mean that the euro had appreciated against the $

if Big Macs were a durable good that could be costly transported between countries, which of the following would present an arbitrage an arbitrage opportunity? check all that apply?

Exporting Big Macs from the Euro area to United State